# Spacetime games subsume causal contextuality scenarios

### v3

@@ 4.6.2 Alternating spacetime game associated with a causal contextuality scenario
We now give the mapping in the opposite direction, i.e., how to associate a spacetime game with a causal contextuality scenario.
Definition 8 (Game associated with a causal contextuality scenario).
@@
Given a causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, we define the associated spacetime game with imperfect information obtained like so:
@@
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(t,x,c)|t\vdash x,c\in\mathcal{C}_{t}\}$
@@
  • $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$
@@
The spacetime game associated with a causal contextuality scenario is alternating.
@@
The game is valid by construction. Thus, we need to show that it is alternating.
2-PLAYERS is true by construction since $\rho$ can only take two values.
BIPARTITE holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
EVEN holds by construction: any node $(t,x,c)$ played by Alfred has $t$ as its parent node, so all root notes must be played by Bob. Likewise, given a node $t$ played by Bob, there is at least one enabling relation $t\vdash x$ that it was taken from. Furthermore, $\mathcal{C}_{t}$ is never empty, so as a consequence, $t$ has at least one child node.
BOB-S (Bob’s information sets are singletons) also holds by construction since $\mathcal{I}_{B}=\mathcal{N}_{B}$.
BOB-A (no unused action at Bob’s nodes) is true: given a node $t$ played by Bob and an action at this node, i.e., a context $c$ in (non-empty) $\mathcal{C}_{t}$, $c$ is used on the edge from $t$ to $x$ for any $x\in c$.
BA1 (nodes played by Alfred have exactly one parent) is true by the construction of $\mathcal{R}_{A}$ because the parent of $(t,x,c)$ is necessarily $t$.
For BA2, let us take a node played by Bob (the left-hand side $t$ of some enabling relation), and that is connected to two nodes played by Alfred. These must be of the form $(t,x,c_{1})$ and $(t,y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(t,x,c)=(t,y,c)$.
AB1 requires that all outgoing edges of the nodes $((u,x,c))_{u,c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$ as well as of $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context.
AB2 requires that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand sides of enabling relations $t$ and $s$. Let us further assume that the nodes for $t$ and $s$ have the same predecessors. This means by construction that $t$ and $s$ have the same domains since the predecessors are made of the domain of $t$ plus contexts. Furthermore, if we assume that the labels of the edges from the predecessors are the same for $t$ and $s$, it implies that $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and BA1 is fulfilled.
Thus, the spacetime game associated with a causal contextuality scenario is always alternating. ∎
@@
The mapping from alternating spacetime games to causal contextuality scenarios as defined in Definition 7 is injective.
@@
We prove that if a spacetime game $\mathcal{G}$ is mapped to a given causal contextuality scenario $(X,O,\vdash,\mathcal{C})$ according to Definition 7, then $\mathcal{G}$ is necessarily, up to an isomorphism [^8] We mean here up to renaming the player names, the labels, etc without affecting the semantics of the game., the game associated with the causal contextuality scenario according to Definition 8.
  • The players
    $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$: this follows from property 2-PLAYERS.
  • Alfred’s information sets
    It must be that $\mathcal{I}_{A}=X$ for the original game by construction of the causal contextuality scenario (first bullet point of Definition 7).
  • Bob’s nodes
    The left-hand sides of the enabling relation are, by definition of the mapping from $\mathcal{G}$, exactly those of the form
    $t(P)=\{(\iota(M),\sigma(M,P))|M\smile P\}$
    for each $P$ that is a node played by Bob in $\mathcal{G}$ and that is the parent of some $M\in x$ such that $t\vdash x$.
    Lemma 2 showed that this is a bijection (because of AB2 and EVEN) and that $\mathcal{N}_{B}$ must therefore match the set of the left-hand sides of the enabling relations: $\mathcal{N}_{B}=\{t|\exists x\in X,t\vdash x\}$.
    As a reminder, this yields the pattern shown in Figure 9. It also follows that $\textsf{enabled}(t)=\textsf{successor}(t)$.
  • Bob’s information sets
    $\mathcal{I}_{B}=\mathcal{N}_{B}$ necessarily because of BOB-S.
  • Bob-Alfred edges
    If we consider an information set $x$, the parent of each node in $x$ must be unique according to BA1. Furthermore, if we consider all enabling relations activating $x$, the parents of nodes in $x$ are exactly and all the values of $t$ such that $t\vdash x$ by definition of the enabling relation in Definition 7. It follows that edges from Bob to Alfred are of the form $(t,(t,x,c))$. Thus, $\mathcal{R}_{A}=\{(t,(t,x,c))|t\vdash x\}$ where the allowed values for $c$, which index the nodes in information set $x$ with parent $t$, $((t,x,c))_{c}$, must still be defined.
  • Bob-Alfred labels
    $\forall(t,(t,x,c))\in\mathcal{R}_{A},\sigma(t,(t,x,c))=c$ follows from the renaming convention of the labels from Bob to Alfred. According to this convention, for any node $t$ played by Bob, there are as many outgoing edges from $t$ as there are pairs $(x,d)$ such that $x\in d$, $d\in\chi(t)$, and $\sigma(t,(x,d))=d$. Thus $d$ also indexes the nodes in the information set $x$ and can be identified with $c$, and the edge going from $t$ to node $(t,x,c)$ must be labeled with $\sigma(t,(t,x,c))=c$.
  • Bob’s actions by node
    We now show that $\chi(t)=\mathcal{C}_{t}$. This amounts to showing that $\chi(t)=\{C\cap\textsf{successor}(t)|C\in\mathcal{C}\}\setminus\{\emptyset\}$.
    Because of the naming convention of Bob’s actions, we know that the labels in $\chi(t)$ are sets of information sets that are successors of $t$. Thus $\chi(t)\subseteq\mathcal{P}(\textsf{successor}(t))$,
    Recall that $\mathcal{C}$ is by construction of the causal contextuality scenario a natural cover. Let us take an element $c$ in $\mathcal{C}_{t}$. By definition, $c$ is not the empty set, and $c$ is the intersection of some context $C\in\mathcal{C}$ with the set of successors of $t$. Thus, there is a corresponding complete history $z$ of the original game such that $\forall i\in\textsf{successor}(t),i\in c\iff z(i)\neq\bot$. This means that $c$ contains exactly the information sets that are activated by $t$ in $z$, and because $c$ is not empty, there is at least one.
    Because of the naming convention of Bob’s nodes, it means that $c$ corresponds to the label born by the edges activating these information sets, and thus $c\in\chi(t)$.
    The converse is also true. If we take an action $c\in\chi(t)$, it labels the edges to all the nodes $(t,x,c)$ such that $x\in c$ because of the naming convention, and $c$ is not the empty set because of EVEN and BOB-A. Thus, there exists a complete history assigning $c$ to $t$ (otherwise $t$ would be an unused node, contradicting that the game is clean) and assigning each $x\in c$ to some value, and no value for the other successors of $t$. The active information sets of Alfred in this complete history form a set that is an element $C$ of $\mathcal{C}$, the natural cover, which is a superset of $c$. Thus $c=C\cap\textsf{successor}(t)$, and $c$ is not the empty set. Thus, $c\in\mathcal{C}_{t}$.
    We have thus proven that $\chi(t)=\mathcal{C}_{t}$.
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ with the same parent $t$ and indexed by the same $c\in\chi(t)$ because of BA2.
  • Alfred’s nodes
    We have shown that the nodes in an information set $x$ with parent $t$ (where $t\vdash x$) are indexed by $\chi(t)$, so that $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\chi(t)\}$. Also, we have shown that $\chi(t)=\mathcal{C}_{t}$. Thus $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\mathcal{C}_{t}\}$.
  • Alfred-Bob edges
    Since we have established the pattern in Figure 9, it follows that edges from Alfred to Bob connect each information set in the domain of $t$ to $t$. Thus, the edges from Alfred to Bob are given by $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$. All these edges must be present (for all $u$ and $c$ such that $(u,x,c)$ is a node) because of AB1’s implying that the nodes in an information set must have the same successors.
  • Alfred-Bob labels
    Again since we have established the pattern in Figure 9, the label on an edge from Alfred’s information set $y$ to Bob’s node $t$ must be $t(y)$. It follows that $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t)=t(x)$
  • Bob’s actions
    $\mathcal{A}_{B}=\{\mathcal{C}_{t}|t\in\mathcal{N}_{B}\}$ because the game has no unused actions and we already showed that $\chi(t)=\mathcal{C}_{t}$.
  • Alfred’s actions by node
    $\forall(t,x,c)\in\mathcal{N}_{A},\chi(t,x,c)=O_{x}$. This is true by the construction of the causal contextuality scenario (second bullet point of Definition 7).
  • Alfred’s actions
    : $\mathcal{A}_{A}=\bigcup_{x}O_{x}$. This is true because there are no unused actions.
  • Assignment to Alfred
    We have $\forall n\in\mathcal{N}_{A},\rho(n)=\text{Alfred}$ by the construction of $\mathcal{N}_{A}$.
  • Assignment to Bob
    We have $\forall n\in\mathcal{N}_{B},\rho(n)=\text{Bob}$ by the construction of $\mathcal{N}_{B}$.
  • All nodes
    We must have $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$ because there are only two players according to 2-PLAYERS.
  • All edges
    $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$ because there are only two players according to 2-PLAYERS.
  • All actions
    $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$ because there are only two players according to 2-PLAYERS.
  • All information sets
    $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$ because there are only two players according to 2-PLAYERS.
5 Concrete impact of the mapping on the simplicity of constructs and proofs

### v4

@@ 3.3.2 Essential surjectivity
We now give the mapping in the opposite direction, i.e., how to associate a spacetime game with a scenario.
Definition 18 (Game associated with a scenario).
@@
Given a scenario $\Gamma=(X,O,\vdash,\mathcal{C})\in\text{ob}(\textbf{Scenario})$, we define the associated spacetime game $\mathcal{G}=G(\Gamma)$ obtained like so:
@@
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(t,x,c)|t\vdash x,c\in\mathcal{C}_{t},x\in c\}$
@@
  • $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{support}(t)\}$
@@
The spacetime game associated with a scenario is alternating:
$\forall\Gamma\in\text{ob}(\textbf{Game}),G(\Gamma)\in\text{ob}(\textbf{Scenario})$
@@
The game is valid by construction of its set of histories. Thus, we need to show that it is alternating.
2-PLAYERS This is true by construction since $\rho$ can only take two values.
BIPARTITE This holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
EVEN This holds by construction: any node $(t,x,c)$ played by Alfred has $t$ as its parent node, so all root notes must be played by Bob. Likewise, given a node $t$ played by Bob, by construction there is at least one enabling relation $t\vdash x$ that it was taken from. Since $\mathcal{C}_{t}$ is never empty, $t$ has at least one child node.
BOB-S Bob’s information sets are singletons by construction since $\mathcal{I}_{B}\equiv\mathcal{N}_{B}$.
BOB-A No action at Bob’s nodes is unused: given a node $t$ played by Bob and an action at this node, i.e., a context $c$ in $\mathcal{C}_{t}$, $c$ is used on the edge from $t$ to $x$ for any $x\in c$.
BA1 Nodes played by Alfred have exactly one parent by the construction of $\mathcal{R}_{A}$ because the parent of $(t,x,c)$ is necessarily $t$.
BA2 Let us take a node played by Bob, which must be the left-hand side $t$ of some enabling relation, and which is connected to two different nodes played by Alfred. These must be of the form $(t,x,c_{1})$ and $(t,y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(t,x,c)=(t,y,c)$. Thus, $(t,x,c_{1})$ and $(t,y,c_{2})$ are in different information sets $x\neq y$.
AB1 requires that all outgoing edges of the nodes $((u,x,c))_{u,c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{support}(t)\}$ as well as of $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context $c$ or on the parent $u$.
AB2 requires that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand sides of enabling relations $t$ and $s$. Let us further assume (reductio ad absurdum) that the nodes $t$ and $s$ have the same predecessors and the corresponding edges have the same labels. This means by construction of $\mathcal{R}_{B}$ that $t$ and $s$ have the same domains. Furthermore, $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and BA2 is fulfilled.
Thus, the spacetime game associated with a scenario is always alternating. ∎
@@
The functor $F:\textbf{Game}\rightarrow\textbf{Scenario}$ is essentially surjective.
@@
[…]
Let us also denote $\mathcal{G}=G(\Gamma)$.
We now show that $\Gamma^{\prime}$ and $\Gamma$ are isomorphic.
First, by construction, $X=\mathcal{I}_{A}\equiv X^{\prime}$ so we can identify $X$ and $X^{\prime}$ and we can define $\pi^{\prime}=\pi=\text{Id}_{X}$.
By construction, all local cover restrictions of $\mathcal{C}$ and $\mathcal{C}$ are the same and they are causally secured. Thus, they are both maximal and both identical. It follows that $\pi^{\prime}$ entails a simplicial map from $\mathcal{C}$ to $\mathcal{C}^{\prime}$ and that $\pi$ entails a simplicial map from $\mathcal{C}^{\prime}$ to $\mathcal{C}$.
Second, for any $x^{\prime}\in X^{\prime}=X$, $O^{\prime}_{x^{\prime}}=\chi(x^{\prime})=O_{x^{\prime}}$ so we can define $\alpha_{x}=\alpha_{x}^{\prime}=\text{Id}_{O_{x}}$ for any $x\in X$.
We now consider the scenario morphism $\gamma:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi^{\prime},\alpha)$, and the morphism $\gamma^{\prime}:\Gamma\rightarrow\Gamma^{\prime}$ with the same components $(\pi,\alpha^{\prime})$. Note that $\gamma$ and $\gamma^{\prime}$ have in fact the same components, but swapped domain and codomain.
If we show that $\gamma$ and $\gamma^{\prime}$ are indeed scenario morphisms, and that $\gamma\circ\gamma^{\prime}=\text{Id}_{\Gamma}$ and $\gamma^{\prime}\circ\gamma=\text{Id}_{\Gamma^{\prime}}$, then we have shown that $\Gamma$ and $\Gamma^{\prime}$ are isomorphic.
First, let us remark that by definition of $F$ and $G$, $\Gamma$ and $\Gamma^{\prime}$ have the same set of measurements $X=X^{\prime}$, which correspond to $\mathcal{I}_{A}$ in $\mathcal{G}$.
Let there be $x,y\in X$ such that [^16] We use $\tau^{\prime}$ to clarify that this refers to the enabling relations in $\Gamma^{\prime}$ to avoid any ambiguity, as $\Gamma$ and $\Gamma^{\prime}$ may have different enabling relations a priori.
$\tau^{\prime}(\pi^{\prime}(x))=\tau^{\prime}(\pi^{\prime}(y))$
Since $\pi^{\prime}$ is the identity, it follows that
$\tau^{\prime}(x)=\tau^{\prime}(y)$
By definition of $\Gamma^{\prime}$ as $F(\mathcal{G})$, $\tau^{\prime}(x)$ corresponds to some node $t$ in $\mathcal{G}$ that is a parent of $x$, its support is given by the parent information sets of $t$ in $\Gamma^{\prime}$, and its values $\tau^{\prime}(x)(x_{i})$ by the labels $\sigma(x_{i},t)$ on the edges from each information set parent $x_{i}$ of $t$.
By definition of $\mathcal{G}$ as $G(\Gamma)$, the parent information sets of $t$ in $\mathcal{G}$ are all the measurements in the support of $t$ in $\Gamma$, so the support of $\tau^{\prime}(x)$ is exactly the support of $\tau(x)$. Furthermore, the label between any parent information set $x_{i}$ of $t$ is, which is $\tau^{\prime}(x)(x_{i})$, is also by definition of $\mathcal{G}$, $\tau(x)(x_{i})$. Thus, $\tau^{\prime}(x)$ and $\tau(x)$ also have the same values and are thus identical. This shows that $\tau^{\prime}(x)=\tau(x)$: the enabling relations of $\Gamma$ and $\Gamma^{\prime}$ are the same. With the same argument $\tau^{\prime}(y)=\tau(y)$ and by transitivity, $\tau(x)=\tau(y)$.
This proves the first rule.
The second rule follows directly from the fact that $\pi^{\prime}$ is the identity, thus, any measurement is in the image of $\pi^{\prime}$.
Let there be $x,y\in X$ such that $\pi^{\prime}(y)\in\textsf{support}(\tau^{\prime}(\pi^{\prime}(x)))$ which is to say $y\in\textsf{support}(\tau^{\prime}(x))$.
Then,
$\alpha_{y}\left[\tau^{\prime}(\pi^{\prime}(x))(\pi^{\prime}(y)\right]=\alpha_{y}\left[\tau^{\prime}(x)(y)\right]=\alpha_{y}\left[\tau(x)(y)\right]=\tau(x)(y)$
This fulfills the third constract. Thus, $\gamma$ is a scenario morphism.
The proof that $\gamma^{\prime}$ is a morphism proceeds similarly.
Thus, $\Gamma$ and $\Gamma^{\prime}$ are isomorphic and $F$ is essentially surjective.
3.3.3 Fullness
We now turn to the fullness of the functor. This means that it maps morphisms, for any pair of games, surjectively.
Theorem 21.
The functor $F:\textbf{Game}\rightarrow\textbf{Scenario}$ is full.
Proof.
We need to show that F maps morphisms between two given games surjectively.
Let us consider two games $\mathcal{G}\in\textsf{ob}(\textbf{Game})$ and $\mathcal{G^{\prime}}\in\textsf{ob}(\textbf{Game})$ and their corresponding scenarios $\Gamma=F(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ and $\Gamma^{\prime}=F(\mathcal{G^{\prime}})=(X^{\prime},O^{\prime},\vdash^{\prime},\mathcal{C^{\prime}})$. Let us consider a morphism $\mu:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi^{\prime},\alpha)$. We need to show that there exists some $\gamma:\mathcal{G^{\prime}}\rightarrow\mathcal{G}$ such that $F(\gamma)=\mu$.
By definition of $F$, $X=\mathcal{I}_{A}$, $O_{x}=\chi(x)$ for every $x\in X$, $X^{\prime}=\mathcal{I}_{A}^{\prime}$, $O_{x^{\prime}}=\chi^{\prime}(x^{\prime})$ for every $x^{\prime}\in X^{\prime}$. Because of these equalities, we can soundly define the second component of our game morphism as $\beta=\alpha$.
$\pi^{\prime}$ naturally maps $\mathcal{I}_{A}$ to $\mathcal{I}_{A}^{\prime}$, but we need to extend it to map $\mathcal{N}_{A}$ to $\mathcal{N}_{A}^{\prime}$ in order to define $\nu^{\prime}$.
Because of Lemma 14, there is a bijection between the left-hand-sides of the enabling relations in $\Gamma$ and $\mathcal{N}_{B}$. There is also a bijection between the left-hand-sides of the enabling relations in $\Gamma^{\prime}$ and $\mathcal{N^{\prime}}_{B}$. The definition of $F$ also implies that $\tau$ is identified with parent.
Let $n\in\mathcal{N}_{A}$, and let us define $t=\text{parent}(n)$, $x=\iota(n)$, $x^{\prime}=\pi^{\prime}(x)=\iota(\pi^{\prime}(n))$, $t^{\prime}=\tau(x^{\prime})$.
Recall that the structure of the edges and children below $t$ structurally form a simplicial complex whose vertices are $t$’s children information sets, and whose facets are the actions in $\chi(t$). Likewise the structure of the edges and children below $t^{\prime}$ structurally form a simplicial complex whose vertices are $t^{\prime}$’s children information sets, and whose facets are the actions in $\chi^{\prime}(t^{\prime})$.
If we apply $\pi^{\prime}$ to the vertices of the simplicial complex at $t$, it can be extended to a simplicial map that in particular maps the facets in $\chi(t)$ to the facets in $\chi^{\prime}(t^{\prime})$.
Thus $\sigma(t,n)$ is mapped to some element $\pi^{\prime}(\sigma(t,n))\in\chi^{\prime}(t^{\prime})$ and since $n\in\sigma(t,n)$ we have $x^{\prime}=\pi^{\prime}(n)\in\pi^{\prime}(\sigma(t,n))$ by definition of a simplicial map.
Since $\pi^{\prime}(\sigma(t,n))\in\chi^{\prime}(t^{\prime})$ and because of the structure of the simplicial complex at $t^{\prime}$, $\pi^{\prime}(\sigma(t,n))$ is a facet thereof, and there is exactly one node $n^{\prime}\in x^{\prime}$ such that $t^{\prime}\smile n^{\prime}$ and $\sigma^{\prime}(t^{\prime},n^{\prime})=\pi^{\prime}(\sigma(t,n))$. We define $\nu^{\prime}(n)=n^{\prime}$.
The morphism $\gamma=(\nu^{\prime},\alpha)$ is indeed a game morphism:
  • The universe nodes are mapped by $\nu^{\prime}$ in a way consistent with the information sets by construction since it maps them according to $\pi^{\prime}$.
  • Because $\mu$ is a scenario morphism, $\pi^{\prime}$ preserves the causal bridge, and thus:
    $\forall x,y\in X,\tau^{\prime}(\pi^{\prime}(x))=\tau^{\prime}(\pi^{\prime}(y))\implies\tau(x)=\tau(y)$
    Since $\tau$ is identified with parent by functor $F$, this implies:
    $\forall x,y\in X,\textsf{parent}(\nu^{\prime}(x))=\textsf{parent}(\nu^{\prime}(y))\implies\textsf{parent}(x)=\textsf{parent}(y)$
    If we now take $n,m\in\mathcal{N}_{A}$ and we have $\textsf{parent}(\nu^{\prime}(m))=\textsf{parent}(\nu^{\prime}(n))$, this implies that $\textsf{parent}(\nu^{\prime}(\iota(m)))=\textsf{parent}(\nu^{\prime}(\iota(n)))$ which implies $\textsf{parent}(\iota(m))=\textsf{parent}(\iota(n))$ which implies $\textsf{parent}(m)=\textsf{parent}(n)$.
    This also defines $\nu$ as an implied morphism component.
  • Let n be a universe node in $\mathcal{G}$ with $\iota(n)=y$ and $t^{\prime}\in\mathcal{D}_{\nu}$ an observer node in $\mathcal{G}$ such that $\nu^{\prime}(n)\smile t^{\prime}$. Since $t^{\prime}\in\mathcal{D}_{\nu}$ there exists some $x\in X$ such that $\textsf{parent}(\nu^{\prime}(x))=t^{\prime}$. This in turn implies $\tau(\pi^{\prime}(x))=\hat{t}^{\prime}$
    Since $\nu^{\prime}(n)\smile t^{\prime}$ we also have $\pi^{\prime}(y)\smile t^{\prime}$ and thus $\pi^{\prime}(y)\in\textsf{support}(\hat{t^{\prime}})=\textsf{support}(\tau(\pi^{\prime}(x)))$.
    Because $\mu$ is a scenario morphism, we have successively:
    $\alpha_{y}\left[\tau(\pi^{\prime}(x))(\pi^{\prime}(y)\right]=\tau(x)(y)$
    $\beta_{y}\left[\tau(\pi^{\prime}(x))(\pi^{\prime}(y)\right]=\tau(x)(y)$
    $\beta_{y}\left[\sigma(\pi^{\prime}(y),\textsf{parent}(\pi^{\prime}(x))\right]=\sigma(y,\textsf{parent}(x))$
    $\beta_{y}\left[\sigma(\pi^{\prime}(y),t^{\prime})\right]=\sigma(y,\nu(t))$
    $\beta_{y}\left[\sigma(\pi^{\prime}(n),t^{\prime})\right]=\sigma(n,\nu(t))$
    which proves the third rule.
  • Let t’ be an observer node in $\mathcal{D}_{\nu}$. Let n be a universe node in $\mathcal{G}$ such that $\nu(t^{\prime})\smile n$. $\pi^{\prime}$ restricted to the children of $\nu(t^{\prime})$ forms a simplicial map from $\chi(\nu(t^{\prime}))$ into $\chi(t^{\prime})$ and by definition of $\nu^{\prime}(n)$:
    $\pi^{\prime}(\sigma(\nu(t^{\prime}),n))=\sigma^{\prime}(t^{\prime},\nu^{\prime}(n))$
    Seeing $\nu^{\prime}$ as a simplicial map, we also have
    $\nu^{\prime}(\sigma(\nu(t^{\prime}),n))=\sigma^{\prime}(t^{\prime},\nu^{\prime}(n))$
    which is the required equality.
  • Let $x^{\prime}$ be a universe information set in $\mathcal{G^{\prime}}$ (i.e., a measurement in X’) and $t^{\prime}$ an observer node in $\mathcal{G^{\prime}}$ such that $x^{\prime}\smile t^{\prime}$ with $t^{\prime}$ corresponds to the left hand side of some enabling relation $t^{\prime}\vdash\nu^{\prime}(x)=\pi^{\prime}(x)$ in $\Gamma^{\prime}$. It implies that $x^{\prime}\in\textsf{support}(\tau(\pi^{\prime}(x))$. Since $\mu$ is a morphism, then there exists some $x\in X=\mathcal{I}_{A}$ so that $\pi^{\prime}(x)=\nu^{\prime}(x)=x^{\prime}$. This proves rule 5.
Thus, $\gamma$ is a morphism and $F(\gamma)=\mu$. Thus, F maps morphisms surjectively. ∎
3.3.4 Faithfulness
We finally turn to the faithfulness of the functor. This means that it maps morphisms, for any pair of games, injectively.
Theorem 22.
The functor $F:\textbf{Game}\rightarrow\textbf{Scenario}$ is faithful.
Proof.
We need to show that F maps morphisms between two given games injectively.
Let us consider two games $\mathcal{G}\in\textsf{ob}(\textbf{Game})$ and $\mathcal{G^{\prime}}\in\textsf{ob}(\textbf{Game})$ and their corresponding scenarios $\Gamma=F(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ and $\Gamma^{\prime}=F(\mathcal{G^{\prime}})=(X^{\prime},O^{\prime},\vdash^{\prime},\mathcal{C^{\prime}})$.
Let us consider two morphisms $\gamma_{1}:\mathcal{G^{\prime}}\rightarrow\mathcal{G}$ with components $(\nu_{1}^{\prime},\beta_{1})$ (and implicitly $\nu_{1}$) and $\gamma_{2}:\mathcal{G^{\prime}}\rightarrow\mathcal{G}$ with components $(\nu_{2}^{\prime},\beta_{2})$ (and implicitly $\nu_{2}$) such that $F(\gamma_{1})=f(\gamma_{2})=\mu$. We now want to prove that $\gamma_{1}=\gamma_{2}$.
Let us denote the components of $\mu$ as $(\pi^{\prime},\alpha)$.
First, by definition of $F$, $\nu_{1}^{\prime}$ and $\nu_{2}^{\prime}$ when considered as functions on information sets are both equal to $\pi^{\prime}$ and thus coincide on information sets. In particular:
$\{\textsf{parent}(\nu_{1}^{\prime}(x))|x\in\mathcal{I}_{A}\}=\{\textsf{parent}(\nu_{2}^{\prime}(x))|x\in\mathcal{I}_{A}\}$
which implies:
$\{\textsf{parent}(\nu_{1}^{\prime}(n))|n\in\mathcal{N}_{A}\}=\{\textsf{parent}(\nu_{2}^{\prime}(n))|n\in\mathcal{N}_{A}\}$
Thus, $\nu_{1}$ and $\nu_{2}$ have the same domain of definition.
Let $x\in\mathcal{I}_{A}$. Then because $\gamma_{1}$ is a morphism, we have
$\nu_{1}(\text{parent}(\nu^{\prime}_{1}(x)))=\text{parent}(x)$
And thus, by definition of F:
$\nu_{1}(\text{parent}(\pi^{\prime}(x)))=\text{parent}(x)$
The same reasoning for $\gamma_{2}$, which is also a morphism, leads to:
$\nu_{2}(\text{parent}(\nu^{\prime}_{2}(x)))=\text{parent}(x)$
And thus, by definition of F:
$\nu_{2}(\text{parent}(\pi^{\prime}(x)))=\text{parent}(x)$
Thus
$\forall x\in\mathcal{I}_{A},\nu_{1}(\text{parent}(\pi^{\prime}(x)))=\nu_{2}(\text{parent}(\pi^{\prime}(x)))$
This means that $\nu_{1}=\nu_{2}$. We denote $\nu=\nu_{1}=\nu_{2}$ in the remainder of this proof.
Let $x\in X=\mathcal{I}_{A}$. By definition of F applied to $\gamma_{1}$, we have $\alpha_{x}(o^{\prime})=\beta_{1,x}(o^{\prime})$ for any o’ in $\mathcal{O^{\prime}}_{\pi^{\prime}(x)}$. By definition of F applied to $\gamma_{2}$, we have $\alpha_{x}(o^{\prime})=\beta_{x,2}(o^{\prime})$ for any o’ in $\mathcal{O}_{\pi^{\prime}(x)}$. Thus $\beta_{1,x}(o^{\prime})=\beta_{2,x}(o^{\prime})$ for any $o^{\prime}\in\mathcal{O^{\prime}}_{\pi^{\prime}(x)}$ that is for any $o^{\prime}\in\chi^{\prime}(\pi^{\prime}(x))$. Thus, $\beta_{1}$ and $\beta_{2}$ coincide for all $x$ and $o^{\prime}$:
$\beta_{1}=\beta_{2}$
Consider some $t^{\prime}$ in $\mathcal{N}_{B}^{\prime}$. We saw that the labels in $\chi^{\prime}(t^{\prime})$ form a simplicial complex whose vertices are the information sets that are the children of $t^{\prime}$. To each face $a^{\prime}=\chi^{\prime}(t^{\prime})$ and each vertex $x^{\prime}\in a^{\prime}$ corresponds exactly one node $n^{\prime}\in x^{\prime}$ such that $\sigma^{\prime}(t^{\prime},n^{\prime})=a^{\prime}$.
If we now consider $t=\nu(t^{\prime})$ in $\mathcal{N}_{B}$, we also have that the labels in $\chi(t)$ form a simplicial complex whose vertices are the information sets that are the children of $t$. To each face $a=\chi(t)$ and each vertex $x\in a$ corresponds exactly one node $n\in x$ such that $\sigma(t,n)=a$.
Furthermore, because $\gamma_{1}$ is a morphism, $\nu_{1}^{\prime}$ seen as a map between information sets forms a simplicial map from $\chi(t)$ to $\chi^{\prime}(t^{\prime})$. Likewise, because $\gamma_{1}$ is a morphism, $\nu_{2}^{\prime}$ seen as a map between information sets forms a simplicial map from $\chi(t)$ to $\chi^{\prime}(t^{\prime})$.
Since $\nu^{\prime}_{1}$ and $\nu_{2}^{\prime}$ coincide on all vertices (which are the information sets), it follows that $\nu^{\prime}_{1}$ and $\nu^{\prime}_{2}$ (seen as simplicial maps $\chi(t)$ to $\chi^{\prime}(t^{\prime})$) coincide fully, i.e., they map the faces (actions) in the same way. Consequently, they also map the nodes (which are face-vertex pairs) in the same way. Thus, $\nu_{1}^{\prime}$ and $\nu_{2}^{\prime}$ coincide on $\mathcal{N}_{A}$.
We can now conclude with the proof of Theorem 11.
Proof.
F is a functor from Game to Scenario that is essentially surjective and fully faithful. It is thus a witness to the equivalence of Game and Scenario. ∎
4 Discussion